# Reinforcement Learning Enhanced LLMs: A Survey

### v2

@@ Abstract
This paper surveys research in the rapidly growing field of enhancing large language models (LLMs) with reinforcement learning (RL), a technique that enables LLMs to improve their performance by receiving feedback in the form of rewards based on the quality of their outputs, allowing them to generate more accurate, coherent, and contextually appropriate responses. In this work, we make a systematic review of the most up-to-date state of knowledge on RL-enhanced LLMs, attempting to consolidate and analyze the rapidly growing research in this field, helping researchers understand the current challenges and advancements. Specifically, we (1) detail the basics of RL; (2) introduce popular RL-enhanced LLMs; (3) review researches on two widely-used reward model-based RL techniques: Reinforcement Learning from Human Feedback (RLHF) and Reinforcement Learning from AI Feedback (RLAIF); and (4) explore Direct Preference Optimization (DPO), a set of methods that bypass the reward model to directly use human preference data for aligning LLM outputs with human expectations. We will also point out current challenges and deficiencies of existing methods and suggest some avenues for further improvements. Project page of this work can be found at our latest repo.
footnotetext: The University of Melbourne, Zhejiang University, CFAR and IHPC, A*STAR, Singapore, Nanyang Technological University, University of Washingtonfootnotetext: Email: shuhewang@student.unimelb.edu.aufootnotetext: Project page of this work can be found at: https://github.com/ShuheWang1998/Reinforcement-Learning-Enhanced-LLMs-A-Surveyfootnotetext: * The latest update was on Dec. 3, 2024 (Version 1).
@@ 1 Introduction
Large language models Jiang et al. (2023); OpenAI (2023); Dubey et al. (2024) are sophisticated language models pre-trained on extensive text data, allowing them to produce coherent and fluent responses to diverse inputs. However, the interaction capabilities of these pre-trained LLMs can be inconsistent, sometimes leading to responses that, while technically correct, may be harmful, biased, misleading, or irrelevant to users’ needs. Therefore, it is crucial to align the outputs of pre-trained LLMs with human preferences before they can be effectively applied to various natural language tasks (Wang et al., 2023b; Wan et al., 2023; Sun et al., 2023c, b; Giray, 2023; Zhang, 2023; Long, 2023; Sun, 2023; Gao et al., 2023; Paranjape et al., 2023; Sun et al., 2023a; Diao et al., 2023; Wang et al., 2023a; Zhang et al., 2023b; Sun et al., 2023d; Liu et al., 2024d; Yao et al., 2024; Liu et al., 2024c; Lee et al., 2024; Kambhampati, 2024; Wang et al., 2024c).
Previously, a widely adopted approach for aligning the outputs of pre-trained LLMs with human preferences has been supervised fine-tuning (SFT) (Hu et al., 2021; Mishra et al., 2021; Wang et al., 2022; Du et al., 2022; Dettmers et al., 2023; Taori et al., 2023; Zhang et al., 2023a; Chiang et al., 2023; Xu et al., 2023; Peng et al., 2023; Mukherjee et al., 2023; Li et al., 2023; Ding et al., 2023; Luo et al., 2023; Wang et al., 2024d; Zhou et al., 2024). This method further trains LLMs on (Instruction, Answer) pairs, where "Instruction" represents the human prompt given to the model, and "Answer" is the target output that follows the instruction. SFT helps guide LLMs to produce responses that adhere to specific characteristics or domain knowledge, making it possible for humans to interactive with LLMs. Despite its effectiveness, SFT has limitations: during training, the model is constrained to learn specific answers we provide, with metrics like perplexity (PPL) penalizing synonym use. On one hand, this can hinder the LLM’s ability to generalize, as tasks like writing and summarization have multiple valid phrasings. On the other hand, it may cause poor performance in aligning with human preferences, as no direct human feedback is incorporated into the training process.
To alleviate the above issues, reinforcement learning (RL) is adopted in aligning the outputs of LLMs with human preferences, which can be decomposed into three steps: (1) First, before fine-tuning, a reward model (or reward function) is trained to approximate human preferences and score different LLM outputs; (2) Then, during each fine-tuning iteration, given a single instruction, the LLM generates multiple responses, each of which is scored by the trained reward model; (3) Finally, policy optimization, an RL optimization technique, updates the LLM’s weights to improve predictions based on these preference scores. Fine-tuing LLMs with RL tackles the aforementioned issues simultaneously. In one line, rather than being restricted to learning a specific answer, RL adjusts the LLM based on various preference scores, rewarding any valid, well-phrased responses. In the other line, the reward model is designed to approximate human preferences, enabling direct training on human preferences and fostering the LLM’s capacity for impressive creativity.
@@
In this paper, we organize the most up-to-date state of knowledge on reinforcement learning (RL) in large language models (LLMs), attempting to consolidate and analyze the rapidly growing research in this field, helping researchers understand the current landscape, challenges, and advancements. Specifically,
@@ 2.2 RL for LLMs
RL Enhanced LLMs Organization # Params RL Methods Instruct-GPT (Ouyang et al., 2022) [Uncaptioned image] 1.3B, 6B, 175B RLHF, PPO GPT-4 (OpenAI, 2023) [Uncaptioned image] - RLHF, PPO, RBRM Gemini (Team et al., 2023) [Uncaptioned image] - RLHF InternLM2 (Cai et al., 2024) [Uncaptioned image] 1.8B, 7B, 20B RLHF, PPO Claude 3 (Anthropic, 2024) [Uncaptioned image] - RLAIF Reka (Team et al., 2024c) [Uncaptioned image] 7B, 21B RLHF, PPO Zephyr (HuggingFaceH4, 2024) [Uncaptioned image] 141B-A39B ORPO Phi-3 (Abdin et al., 2024) [Uncaptioned image] 3.8B, 7B, 14B DPO DeepSeek-V2 (Liu et al., 2024a) [Uncaptioned image] 236B-A21B GRPO ChatGLM (GLM et al., 2024) [Uncaptioned image] 6B, 9B ChatGLM-RLHF Nemotron-4 340B (Adler et al., 2024) [Uncaptioned image] 340B DPO, RPO Llama 3 (Dubey et al., 2024) [Uncaptioned image] 8B, 70B, 405B DPO Qwen2 (Yang et al., 2024a) [Uncaptioned image] (0.5-72)B, 57B-A14B DPO Gemma2 (Team et al., 2024b) [Uncaptioned image] 2B, 9B, 27B RLHF Starling-7B (Zhu et al., 2024) [Uncaptioned image] 7B RLAIF, PPO Athene-70B (Nexusflow, 2024) [Uncaptioned image] 70B RLHF Hermes 3 (Teknium et al., 2024) [Uncaptioned image] 8B, 70B, 405B DPO o1 (OpenAI, 2024b) [Uncaptioned image] - RL through CoT

### v3

@@ Abstract
Reinforcement learning (RL) enhanced large language models (LLMs), particularly exemplified by DeepSeek-R1, have exhibited outstanding performance. Depsite the effectiveness in improving LLM capabilities, its implementation remains highly complex, requiring complex algorithms, reward modeling strategies, and optimization techniques. This complexity poses challenges for researchers and practitioners in developing a systematic understanding of RL-enhanced LLMs. Moreover, the absence of a comprehensive survey summarizing existing research on RL-enhanced LLMs has limited progress in this domain, hindering further advancements.
In this work, we are going to make a systematic review of the most up-to-date state of knowledge on RL-enhanced LLMs, attempting to consolidate and analyze the rapidly growing research in this field, helping researchers understand the current challenges and advancements. Specifically, we (1) detail the basics of RL; (2) introduce popular RL-enhanced LLMs; (3) review researches on two widely-used reward model-based RL techniques: Reinforcement Learning from Human Feedback (RLHF) and Reinforcement Learning from AI Feedback (RLAIF); and (4) explore Direct Preference Optimization (DPO), a set of methods that bypass the reward model to directly use human preference data for aligning LLM outputs with human expectations. We will also point out current challenges and deficiencies of existing methods and suggest some avenues for further improvements. Project page of this work can be found at our latest repo.
footnotetext: The University of Melbourne, Zhejiang University, CFAR and IHPC, A*STAR, Singapore, Nanyang Technological University, University of Washingtonfootnotetext: Email: shuhewang@student.unimelb.edu.aufootnotetext: Project page of this work can be found at: https://github.com/ShuheWang1998/Reinforcement-Learning-Enhanced-LLMs-A-Surveyfootnotetext: * The latest update was on Feb. 24, 2025 (Version 2).
@@ 1 Introduction
Recently, represented by DeepSeek-R1 DeepSeek-AI et al. (2025), reinforcement learning (RL) enhanced large language models (LLMs) OpenAI (2023, 2024a); Team et al. (2024b); GLM et al. (2024); Adler et al. (2024); Yang et al. (2024a); AI et al. (2024); Team et al. (2025), have demonstrated remarkable performance (Wang et al., 2023b; Wan et al., 2023; Sun et al., 2023c, b; Giray, 2023; Zhang, 2023; Long, 2023; Sun, 2023; Gao et al., 2023; Paranjape et al., 2023; Sun et al., 2023a; Diao et al., 2023; Wang et al., 2023a; Zhang et al., 2023b; Sun et al., 2023d; Liu et al., 2024d; Yao et al., 2024; Liu et al., 2024c; Lee et al., 2024; Kambhampati, 2024; Wang et al., 2024c), attracting widespread attention from the research community. A key factor contributing to this success is the integration of reinforcement learning techniques, which have proven to be a powerful approach to enhancing LLM capabilities.
The process of training LLMs using RL can be divided into three main steps: (1) Reward Model Training: before fine-tuning, a reward model (or reward function) is trained to approximate human preferences and evaluate different LLM outputs; (2) Preference-Based Fine-Tuning: during each fine-tuning iteration, the LLM generates multiple responses for a given instruction, each of which is scored using the trained reward model; and (3) Policy Optimization: using RL optimization techniques, the model’s weights are updated based on preference scores, improving response generation. Incorporating reinforcement learning into LLMs allows models to adjust dynamically based on varied preference scores, rather than being confined to a single predetermined answer. This enables them to produce well-structured, contextually appropriate responses. Furthermore, RL enables direct training on human preferences, enhancing the LLM’s ability to generate creative, high-quality outputs that align more closely with human expectations.
Although reinforcement learning (RL) is highly effective in enhancing LLM capabilities, its implementation remains highly complex, requiring sophisticated algorithms, reward modeling strategies, and optimization techniques. This complexity poses challenges for researchers and practitioners in developing a systematic understanding of RL-enhanced LLMs. Moreover, the absence of a comprehensive survey summarizing existing research on RL-enhanced LLMs has limited progress in this domain, hindering further advancements and making it difficult to consolidate knowledge and drive innovation in the field.
To address this issue, in this paper, we provide a systematic review of both the challenges and opportunities involved in reinforcement learning-enhanced LLMs. Specifically, we aim to identify the key obstacles in applying reinforcement learning to LLMs and offer a comprehensive analysis of recent efforts to overcome them:
(1) At the reward model training level, we analyze the evolution of reward models, including Reinforcement Learning from Human Feedback (RLHF) and Reinforcement Learning from AI Feedback (RLAIF). Our discussion covers their effectiveness, limitations, evaluation methods, and challenges, such as bias in human annotations, generation of out-of-distribution content, and issues related to human interpretability.
(2) At preference fine-tuning level, we explore approaches for aligning LLMs with human preferences, including direct preference optimization (DPO) and its variations. Additionally, we examine the impact of preference data collection, different optimization functions, and strategies designed to maintain safety throughout the alignment process.
@@
By presenting a comprehensive overview of RL-enhanced LLMs, this survey aims to bridge the knowledge gap in the field, attempting to consolidate and analyze the rapidly growing research in this field, helping researchers understand the current landscape, challenges, and advancements. The rest of this survey is organized as follows:
@@ 2.2 RL for LLMs
RL Enhanced LLMs Organization # Params RL Methods DeepSeek-R1 (DeepSeek-AI et al., 2025) [Uncaptioned image] 671B-A31B RL through CoT Kimi-k1.5 (Team et al., 2025) [Uncaptioned image] - RL through CoT o1 (OpenAI, 2024b) [Uncaptioned image] - RL through CoT Hermes 3 (Teknium et al., 2024) [Uncaptioned image] 8B, 70B, 405B DPO Athene-70B (Nexusflow, 2024) [Uncaptioned image] 70B RLHF Starling-7B (Zhu et al., 2024) [Uncaptioned image] 7B RLAIF, PPO Gemma2 (Team et al., 2024b) [Uncaptioned image] 2B, 9B, 27B RLHF Qwen2 (Yang et al., 2024a) [Uncaptioned image] (0.5-72)B, 57B-A14B DPO Llama 3 (Dubey et al., 2024) [Uncaptioned image] 8B, 70B, 405B DPO Nemotron-4 340B (Adler et al., 2024) [Uncaptioned image] 340B DPO, RPO ChatGLM (GLM et al., 2024) [Uncaptioned image] 6B, 9B ChatGLM-RLHF DeepSeek-V2 (Liu et al., 2024a) [Uncaptioned image] 236B-A21B GRPO Phi-3 (Abdin et al., 2024) [Uncaptioned image] 3.8B, 7B, 14B DPO Zephyr (HuggingFaceH4, 2024) [Uncaptioned image] 141B-A39B ORPO Reka (Team et al., 2024c) [Uncaptioned image] 7B, 21B RLHF, PPO Claude 3 (Anthropic, 2024) [Uncaptioned image] - RLAIF InternLM2 (Cai et al., 2024) [Uncaptioned image] 1.8B, 7B, 20B RLHF, PPO Gemini (Team et al., 2023) [Uncaptioned image] - RLHF GPT-4 (OpenAI, 2023) [Uncaptioned image] - RLHF, PPO, RBRM Instruct-GPT (Ouyang et al., 2022) [Uncaptioned image] 1.3B, 6B, 175B RLHF, PPO
@@ 3 Popular LLMs Enhanced by RL
3.1 DeepSeek-R1
DeepSeek-R1 (DeepSeek-AI et al., 2025), developed by DeepSeek, is a state-of-the-art reasoning model that achieves performance comparable to OpenAI’s o1 OpenAI (2024b) series models. This work pioneers the application of pure reinforcement learning (RL) to enhance language model reasoning, emphasizing self-evolution rather than relying solely on supervised training data.
During the training stage, DeepSeek-R1 undergoes an alternating training process that integrates supervised fine-tuning (SFT) Zhang et al. (2023a) and reinforcement learning (RL) across four key stages: (1) Initial Cold Start SFT, begining with the collection of thousands of high-quality, readability-focused long-CoT datasets to fine-tune DeepSeek-V3-Base, establishing a strong foundation for subsequent RL training; (2) First Reasoning-Oriented RL Stage, using the methodology named large-scale reasoning-focused RL to further fine-tune the model, and the resulting checkpoint is used to generate additional SFT data for the next training phase; (3) Second SFT Stage, refining the model further by incorporating both reasoning and non-reasoning data. For the reasoning data, it is generated using rejection sampling from the RL checkpoint of the previous stage. For the non-reasoning data, the DeepSeek-V3 pipeline is used to integrate data for writing, factual QA, self-cognition, and translation, including parts of the DeepSeek-V3 SFT dataset; (4) Second RL Stage for All Scenarios, leverating the final RL stage to further align the model with human preferences, improving its helpfulness, harmlessness, and reasoning abilities.
DeepSeek-R1 is an open-source model, made publicly available to the research community to support further advancements in the field. The release process first introduces DeepSeek-R1-Zero, a model derived from DeepSeek-V3-Base, trained using large-scale RL without supervised fine-tuning (SFT). This initial model demonstrates remarkable reasoning improvements, with its pass@1 score on AIME 2024 increasing from 15.6% to 71.0%. With majority voting, the score further rises to 86.7%, matching OpenAI-o1-0912. To address issues such as poor readability and language mixing, and to further improve reasoning capabilities, the authors introduce DeepSeek-R1. This enhanced model integrates a small amount of cold-start data and follows a multi-stage training pipeline, achieving performance on par with OpenAI-o1-1217.
3.2 Kimi-k1.5
Kimi-k1.5 (Team et al., 2025), developed by Moonshot AI, is a multi-modal large language model (LLM) that marks a major breakthrough in scaling reinforcement learning (RL). The model introduces a novel approach by emphasizing long-context scaling, extending the RL context window to 128k, and refining policy optimization techniques.
For training, Kimi-k1.5 undergoes a four-stage training process consisting of: (1) Pre-training, (2) Vanilla supervised fine-tuning (SFT), (3) Long-CoT supervised fine-tuning, and (4) Reinforcement learning (RL). The key innovation lies in the reinforcement learning (RL) phase, where the authors develop a high-quality RL prompt set to guide the model toward robust reasoning while mitigating potential risks such as reward hacking and overfitting to superficial patterns. This prompt set is designed with three essential properties: (1) Diverse coverage, ensuring exposure to a broad range of reasoning challenges; (2) Balanced difficulty, providing a mix of easy, moderate, and complex reasoning tasks; and (3) Accurate evaluability, allowing precise measurement of reasoning performance.
While long-CoT (Chain-of-Thought) models exhibit strong reasoning performance, they tend to consume more tokens at test time compared to standard short-CoT LLMs. To tackle this long2short challenge, the authors propose four key methods to transfer long-CoT reasoning capabilities to short-CoT models: (1) Model Merging, combining a long-CoT model with a shorter model by averaging their weights; (2) Shortest Rejection Sampling, generating multiple responses using the long-CoT model and selecting the shortest correct response for SFT training; (3) Direct Preference Optimization (DPO), constructing pairwise preference data, where shorter correct solutions are treated as positive samples and longer solutions as negative samples, optimizing the model through DPO training; and (4) Long2Short RL, implementing a two-phase RL training approach, where the model first undergoes standard RL training, and then a length penalty is applied, and the maximum response length is reduced, encouraging more concise reasoning while maintaining performance.
Kimi-k1.5 delivers state-of-the-art reasoning performance across various benchmarks and modalities, rivaling OpenAI’s o1 OpenAI (2024b). Moreover, the introduction of long2short techniques significantly enhances short-CoT models, achieving up to 550% improvement over existing models such as GPT-4o Hurst et al. (2024) and Claude 3.5111https://www.anthropic.com/news/claude-3-5-sonnet.